# GOProteinGNN: Leveraging Protein Knowledge Graphs for Protein Representation Learning

### v1

@@ 3.1. GOProteinGNN
We present the GOProteinGNN architecture (Fig. 2b). GOProteinGNN extends the capabilities of PLMs by incorporating knowledge graph information into the models, using our proposed GKI layer (See Section 3.2). We leveraged the [CLS] token representation, to integrate between the proposed GKI and the PLM in a novel manner. This fusion of language model-based representation learning with knowledge graph processing allows GOProteinGNN to generate enriched protein embeddings that capture essential biological contexts and interactions. The detailed algorithm is presented in Algorithm 1.
@@
3.1.1. Training Phase
In language models, including a specialized classification token ([CLS]) at the beginning of the sequence is a common practice. The final hidden state of this token serves as a comprehensive sequence representation for various classification tasks (Devlin et al., 2019). In our framework, we leverage the [CLS] token representation to integrate external knowledge into our representation framework and enhance the protein representation.
Within the context of a given encoder layer denoted by $l$, we introduce the GKI layer. The GKI operation (See Section 3.2) takes the input representation $[CLS]_{l}$ along with the protein knowledge graph denoted by $G$ (See Section 3.2.1), resulting in a refined representation $[CLS]^{\prime}_{l}$ (line 9):
$[CLS]^{\prime}_{l}=GKI([CLS]_{l},G)\in\mathbb{R}^{D}$
Subsequently, this refined [CLS] token representation assumes the role of the new [CLS] token in the ensuing encoder layer. Then, the model propagates and refines the representations of all other amino acids within the sequence, resulting in a new representation of the amino acid sequence (line 7).
As the protein representation undergoes transformations across the encoder layers, our aim is to uniformly integrate knowledge a total of $k$ times across the $L$ encoder layers. Therefore, we add a GKI layer after every $L\div k$ encoder layer (line 8).
3.1.2. Inference Phase
To eliminate the dependency on a knowledge graph (KG) during inference, we introduce a weight-less injection mechanism. During the inference phase, GOProteinGNN omits the GKI components utilized solely for pre-training. In addition, we incorporate task-specific classification heads to enable predictions for these specific tasks. This separation between pre-training and inference empowers GOProteinGNN to generalize effectively across various protein tasks, including those involving unseen proteins.
3.2. GNN Knowledge Injection (GKI)
GNNs (Scarselli et al., 2009) are powerful neural networks designed for graph-structured data (Liu et al., 2021; Agiollo and Omicini, 2022), making them suitable for learning effectively the protein knowledge graph, similar to common usages in other domains (Sun, 2022; Wang et al., 2021). By utilizing the protein knowledge graph denoted by $G$, we introduce a methodology called GNN Knowledge Injection, denoted by GKI. This method produces an updated protein representation that incorporates information from the knowledge graph into the input representation.
3.2.1. Protein Knowledge Graph
GOProteinGNN is trained on a knowledge graph consisting of two different types of nodes: Proteins and GO-Terms, with Relations that act as edges. Proteins are represented as a sequence of amino acids, while the relation and the GO term are terms of factual knowledge described in natural language. The GO terms cover three aspects of biology: Molecular Function, Cellular Component and Biological Process.
Protein Knowledge Graph Definition
3.2.2. GNN Method
A Protein Knowledge Graph consists of multiple types of relations between proteins and GO terms. In response to the complex diversity of edge types in the graph, we select a GNN that accommodates the multifaceted edge information. Recognizing that conventional GCNs (Kipf and Welling, 2017) lack the capacity to fully exploit this characteristic, we turn to the Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017), a specialized variant of GNN. By doing so, we can effectively harness the rich information present in the knowledge graph to refine and enhance the protein embedding. This approach allows us to leverage the graph structure of the protein knowledge graph and to capture complex relationships and dependencies between proteins and their variable number of functional annotations, leading to more powerful and contextually enriched protein representations, a capability not attainable through other methods, such as embedding concatenation. Moreover, other graph algorithms could also be applied. We compare different graph algorithms in Section 6.1.
3.2.3. Knowledge Injection
Given a protein representation (i.e., the [CLS] representation) $z_{protein}\in\mathbb{R}^{D}$ corresponding to node $i$ in the knowledge graph $G$, set:
$h_{i}^{(0)}=z_{protein}$
The given protein is connected to multiple nodes in the knowledge graph $G$, such as the GO terms nodes. To effectively utilize the textual information of these nodes, the GO term descriptions are processed through a language encoder (See Section 3.2.1) to generate comprehensive knowledge representations. To align these text-based representations with the protein embeddings, we project both types of representations into a unified space by applying an affine transformation.
The $(l+1)$ RGCN layer performs the following update to the node representation $z_{protein}=h_{i}^{(l)}$:
The new protein embedding after the $L$ RGCN layers is given by:
$z^{{}^{\prime}}_{protein}=h_{i}^{(L)}$
where $z^{{}^{\prime}}_{protein}\in\mathbb{R}^{D}$ represents the updated protein embedding that captures essential biological contexts and interactions.
3.3. Masked Language Modeling Objective
$\mathcal{L}_{MLM}=-\sum_{j=0}^{M-1}\log P(a_{(j)}|z_{p},\Theta)$
Table 1. Results on amino acid contact prediction task. Seq indicates the distance (i.e., the number of amino acids) between two selected amino acids. P@L, P@L/2, and P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively. The best result is highlighted in bold.
Method 6 $\leq$ $seq$ $<$ 12 12 $\leq$ $seq$ $<$ 24 24 $\leq$ $seq$
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
LSTM 0.26 0.36 0.49 0.20 0.26 0.34 0.20 0.23 0.27
ResNet 0.25 0.34 0.46 0.28 0.25 0.35 0.10 0.13 0.17
Transformer 0.28 0.35 0.46 0.19 0.25 0.33 0.17 0.20 0.24
ESM-1b (Rives et al., 2019) 0.38 0.48 0.62 0.33 0.43 0.56 0.26 0.34 0.45
KeAP (Zhou et al., 2023) 0.41 0.51 0.63 0.36 0.45 0.54 0.28 0.35 0.43
GOProteinGNN $\textbf{0.43}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.65}^{*}$ $\textbf{0.38}^{*}$ $\textbf{0.49}^{*}$ $\textbf{0.60}^{*}$ $\textbf{0.29}^{*}$ $\textbf{0.38}^{*}$ $\textbf{0.48}^{*}$
4. Empirical Evaluation
In this section we outline our empirical setting. We performed analysis on both intrinsic (e.g., semantic similarity inference of proteins) and extrinsic tasks (e.g., protein-protein interactions), demonstrating the value of the novel representation we provide in real world bioinformatics tasks.
4.1. Pre-Training Dataset
@@
We employed the ProteinKG25 dataset (Zhang et al., 2022) for pre-training, consistent with all knowledge-enhanced baseline methods. ProteinKG25 provides a knowledge graph with aligned descriptions and protein sequences, respectively, to GO terms (Ashburner et al., 2000) and protein entities. GO terms can be a molecule, a cellular component, or a biological process. The GO terms and their relations to proteins are described using natural language. The dataset contains around 5 million triplets, including nearly 600k proteins, 50k attribute terms, and 31 relation terms, with an average of 8.64 relations per protein. We address a data leakage concern by following the methodology proposed by KeAP (Zhou et al., 2023). After removing the leakage from the dataset, their algorithm is considered SOTA across all tasks. Given the robust performance demonstrated by KeAP in various downstream applications, outperforming other methods including OntoProtein (Zhang et al., 2022), we focus our comparative analysis solely on KeAP (Zhou et al., 2023).
4.2. Tasks
We use Tasks Assessing Protein Embeddings (TAPE) (Rao et al., 2019) as the benchmark to evaluate protein representation learning. TAPE is a set of biologically relevant semi-supervised learning tasks spread across different domains of protein biology. All tasks were divided into specific training, validation, and test splits to ensure that each task tested a biologically relevant generalization that transferred to real-life scenarios. Furthermore, we tested our representations on Semantic Similarity Inference and Binding Affinity Estimation tasks from PROBE (Unsal et al., 2022) and on Protein-Protein Interactions (PPI) from GNN-PPI (Lv et al., 2021).
4.2.1. Contact Prediction

### v2

@@ 3. Methods
We present the GOProteinGNN architecture (Fig. 2), which extends the capabilities of PLMs by incorporating knowledge graph information into the models, using our proposed GKI layer. We leveraged the [CLS] token representation, to integrate between the proposed GKI and the PLM in a novel manner. This integration of language model-based representation learning with knowledge graph processing allows GOProteinGNN to generate enriched protein representations that capture essential biological contexts and interactions. The detailed algorithm is presented in Supplementary Material A.
3.1. Protein Knowledge Graph
GOProteinGNN is trained on a knowledge graph consisting of two different types of nodes: Proteins and GO-Terms, with Relations that act as edges. Proteins are represented as a sequence of amino acids, while the relation and the GO term are terms of factual knowledge described in natural language. The GO terms cover three aspects of biology: Molecular Function, Cellular Component, and Biological Process.
Protein Knowledge Graph Definition
3.2. GKI Layer
We introduce a method called GNN Knowledge Injection (GKI), inspired by GNN (Scarselli et al., 2009), a powerful neural network designed for graph-structured data (Liu et al., 2021; Agiollo and Omicini, 2022). GKI is particularly suitable for effectively learning from protein knowledge graphs, analogous to applications in other domains (Sun, 2022; Wang et al., 2021). By utilizing a protein knowledge graph, denoted as $G$, the GKI layer generates an updated protein representation that incorporates information from the knowledge graph into the input representation.
GNN Method
In response to the complex diversity of edge types in the protein knowledge graph, we select a GNN that accommodates the multifaceted edge information. Recognizing that conventional GCNs (Kipf and Welling, 2017) lack the capacity to fully exploit this characteristic, we turn to the Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017), a specialized variant of GNN. By doing so, we can effectively harness the rich information present in the knowledge graph to refine and enhance the protein representation. This approach allows us to leverage the graph structure of the protein knowledge graph and to capture complex relationships and dependencies between proteins and their variable number of functional annotations, leading to more powerful and contextually enriched protein representations, a capability not attainable through other methods, such as representation concatenation. Moreover, other graph algorithms could also be applied. We compare different graph algorithms in Section 6.1.
3.2.1. GKI Procedure
Given a protein representation (i.e., the [CLS] representation) $z_{protein}\in\mathbb{R}^{D}$ corresponding to node $i$ in the knowledge graph $G$, set: $h_{i}^{(0)}=z_{protein}$.
The given protein is connected to multiple nodes in the knowledge graph $G$, such as the GO terms nodes ($V_{terms}$). The $(l+1)$ RGCN layer performs the following update to the node representation $z_{protein}=h_{i}^{(l)}$:
The new protein representation after the $L$ RGCN layers is given by: $z^{{}^{\prime}}_{protein}=h_{i}^{(L)}$ where $z^{{}^{\prime}}_{protein}\in\mathbb{R}^{D}$ represents the updated protein representation that captures essential biological contexts and interactions.
3.3. GOProteinGNN Algorithm
3.3.1. Training Phase
In language models, including a specialized classification token ([CLS]) at the beginning of the sequence is a common practice. The final hidden state of this token serves as a comprehensive sequence representation for various classification tasks (Devlin et al., 2019). In our framework, we uniquely leverage the [CLS] token representation to integrate external knowledge into our representation framework and enhance the protein representation.
Within the context of a given encoder layer denoted by $l$, a GKI layer takes the input representation $[CLS]_{l}$ along with the protein knowledge graph denoted by $G$, resulting in a refined representation $[CLS]^{\prime}_{l}$ (line 9):
$[CLS]^{\prime}_{l}=GKI([CLS]_{l},G)\in\mathbb{R}^{D}$
Subsequently, this refined [CLS] token representation assumes the role of the new [CLS] token in the ensuing encoder layer. Then, using the self-attention (Vaswani et al., 2017) of the encoder layer, the refined representation propagates and refines the representations of all other amino acids within the sequence, resulting in a new representation of the amino acid sequence (line 7).
As the protein representation undergoes transformations across the encoder layers, our aim is to uniformly integrate knowledge a total of $k$ times across the $L$ encoder layers. Therefore, we add a GKI layer after every $L\div k$ encoder layer (line 8). Unlike prior language models that utilize the [CLS] token only in the final output (Devlin et al., 2019), our architecture iteratively integrates knowledge across the encoder layers, ensuring a more consistent and dynamic integration of knowledge throughout the representation process.
Method Short Range Medium Range Long Range
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
LSTM 0.26 0.36 0.49 0.20 0.26 0.34 0.20 0.23 0.27
ResNet 0.25 0.34 0.46 0.28 0.25 0.35 0.10 0.13 0.17
Transformer 0.28 0.35 0.46 0.19 0.25 0.33 0.17 0.20 0.24
ESM-1b (Rives et al., 2019) 0.38 0.48 0.62 0.33 0.43 0.56 0.26 0.34 0.45
ESM-2 (Lin et al., 2023) 0.40 0.50 0.62 0.35 0.44 0.56 0.27 0.35 0.45
KeAP (Zhou et al., 2023) 0.41 0.51 0.63 0.36 0.45 0.54 0.28 0.35 0.43
GOProteinGNN $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{0.39}^{*}$ $\textbf{0.50}^{*}$ $\textbf{0.60}^{*}$ $\textbf{0.30}^{*}$ $\textbf{0.39}^{*}$ $\textbf{0.48}^{*}$
Table 1. Results on amino acid contact prediction task. Short-range, medium-range, and long-range contacts are contacts between positions that are separated by 6 to 11, 12 to 23, and 24 or more positions, respectively. P@L, P@L/2, and P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
3.3.2. Masked Language Modeling Training Objective
$\mathcal{L}_{MLM}=-\sum_{j=0}^{M-1}\log P(a_{(j)}|z_{p},\Theta)$
3.3.3. Inference Phase
During inference, as factual knowledge is not available, and to eliminate its dependency, we introduce a weight-less injection mechanism. Specifically, GOProteinGNN omits the GKI components utilized solely for pre-training. Furthermore, task-specific classification heads are incorporated to enable predictions for these specific tasks. This separation between pre-training and inference allows GOProteinGNN to generalize effectively across various protein tasks, including those involving unseen proteins.
4. Empirical Evaluation
In this section, we outline our empirical setting. We performed extensive analysis on both intrinsic (e.g., semantic similarity inference of proteins) and extrinsic tasks (e.g., protein-protein interactions), demonstrating the value of the novel representation we provide in real-world bioinformatics tasks. Detailed experimental settings, such as implementation details and computational complexity, are provided in the Supplementary Material C. For pre-training, we employed the ProteinKG25 dataset (Zhang et al., 2022), consistent with all knowledge-enhanced baseline methods (See Supplementary Material B for full details). Furthermore, the limitations of knowledge graphs and potential mitigation strategies are discussed in Supplementary Material B.0.1.
4.1. Tasks
We present an extensive empirical evaluation of our method across a wide range of protein-level and amino acid-level representation tasks. We use Tasks Assessing Protein Embeddings (TAPE) (Rao et al., 2019) as a common benchmark to evaluate protein representation learning. TAPE is a set of biologically relevant semi-supervised learning tasks, such as Contact Prediction and Remote Homology Detection, spread across different domains of protein biology. Furthermore, we tested our representations on the Semantic Similarity Inference task from PROBE (Unsal et al., 2022) and on Protein-Protein Interactions (PPI) from GNN-PPI (Lv et al., 2021). All tasks were divided into specific training, validation, and test splits to ensure that each task tested a biologically relevant generalization that transferred to real-life scenarios. See Supplementary Material D for comprehensive task details and Supplementary Material E for an analysis of additional tasks.
4.2. Baselines
Incorporating SOTA models as our baselines, we used ProtBert (Elnaggar et al., 2022) and ESM-1b (Rives et al., 2019) as benchmarks, following KeAP (Zhou et al., 2023) for fair comparison. Furthermore, we added the SOTA ESM-2 (Lin et al., 2023), which improves the ESM-1b model. These models are considered the top-performing models in the absence of factual knowledge integration. Furthermore, we subjected our analysis to the scrutiny of the most recent and advanced knowledge-driven pre-training model, KeAP (Zhou et al., 2023). We exclude the results of OntoProtein (Zhang et al., 2022) in this paper, as it was presented by KeAP (Zhou et al., 2023) that it is significantly inferior on a dataset without leakage (See Supplementary Material B). For each specific task, we followed the convention of selecting SOTA methods, just as prior research has done (See Supplementary Material C.1 for full details on task-specific baselines).
5. Empirical Results
In this section, we present our empirical results. We validate the statistical significance of performance differences, using a one-tailed paired t-test with a 95% confidence level ($p<0.05$), comparing observations from the tested models. Significant results are marked with an asterisk (*) in the tables.
5.1. Contact Prediction
In Table 1, we present GOProteinGNN’s performance on the contact prediction task compared to six baseline methods. GOProteinGNN significantly outperformed all baseline methods across all categories and metrics. In particular, GOProteinGNN showed superior performance compared to ProtBert (Elnaggar et al., 2022) and ESM-2 (Lin et al., 2023), which lack knowledge graph information. This highlights the importance of incorporating knowledge during representation learning, especially for pairwise amino acid tasks (e.g., contact prediction task), where success depends on understanding the proteins’ functions. Moreover, KeAP (Zhou et al., 2023) incorporates a knowledge post-encoder, limiting its capacity to seamlessly integrate external knowledge into protein representations, as reflected in the performance gains of GOProteinGNN over KeAP. These findings highlight the significant impact of incorporating knowledge graph information in GOProteinGNN via the GKI layers, leading to superior performance in the contact prediction task.
5.2. Semantic Similarity Inference
Table 2 presents the results of the semantic similarity inference task. GOProteinGNN outperformed all baselines significantly in capturing semantic similarities, achieving an increase of 23.8% compared to the second-best model. Biological processes play a crucial role in many drug development processes, making this category particularly important as it pertains directly to drug development. The success in this task underscores GOProteinGNN’s efficacy in acquiring enriched representations that encompass the complex biological processes in which proteins are involved, due to the whole protein knowledge graph learning, surpassing the capabilities of alternative approaches. Furthermore, KeAP (Zhou et al., 2023) showed inferior performance compared to ESM-1b (Rives et al., 2019), which does not use factual knowledge. This result is surprising given that both GOProteinGNN and KeAP utilize the same factual knowledge repository, yet KeAP’s method of knowledge injection failed to outperform the model that does not use knowledge at all.
Method Biological Process
MSA Transformer 0.31
ProtT5-XL 0.21
ProtBert 0.35
ESM-1b 0.42
ESM-2 0.41
KeAP 0.41
GOProteinGNN $\textbf{0.52}^{*}$
Table 2. Results on semantic similarity inference task. The Spearman’s rank is the evaluation metric. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
5.3. Protein-Protein Interaction Identification
Table 3 presents the results of the protein-protein interaction (PPI) identification task. GOProteinGNN showed strong performance across the board, achieving the highest F1 scores across all datasets with statistical significance, which can be beneficial for drug discovery or personalized medicine. In particular, GOProteinGNN obtained SOTA results with an increase of approximately 2.1% on the SHS27K dataset, and 2.7% on the SHS148K dataset compared to KeAP (Zhou et al., 2023), with statistical significance.
In conclusion, the results highlight the strengths of our GOProteinGNN, which enables a better biological understanding of the protein, leading to success in predicting interactions between proteins.
Method SHS27K SHS148K
DNN-PPI 48.09 57.40
DPPI 41.43 52.12
PIPR 44.48 61.83
GNN-PPI 63.81 71.37
ProtBert 70.94 70.32
ESM-1b 74.92 77.49
ESM-2 75.05 77.19
KeAP 78.58 77.22
GOProteinGNN $\textbf{80.24}^{*}$ $\textbf{79.27}^{*}$
Table 3. Results on protein-protein interaction (PPI) identification task. The table displays the results in two datasets: SHS27K and SHS148K. The breadth-first search (BFS) method is used. The F1 score is the evaluation metric. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
5.4. Remote Homology Detection
In Table 4, we present GOProteinGNN’s performance compared to seven baseline methods on the remote homology detection task. Despite the challenging nature of the task, due to the large number of potential protein folds, GOProteinGNN significantly outperformed all methods, demonstrating its effectiveness in handling the complexities of remote homology detection. Specifically, GOProteinGNN achieved an advantage over KeAP (Zhou et al., 2023), further affirming its ability to capture essential biological contexts and interactions which is crucial for predicting the protein fold structure, as it processed the whole protein knowledge graph. As remote homology detection is a sequence-level classification task, we can see that both KeAP (Zhou et al., 2023) and ProtBert (Elnaggar et al., 2022) achieved equal results, suggesting KeAP fails to leverage the factual knowledge in the protein level in this task. This task highlights the novelty of our algorithm which allows for the integration of information at both the individual amino acid level and the entire protein level.
Method Homology
LSTM 0.26
ResNet 0.17
Transformer 0.21
ProtBert 0.29
ProteinBert 0.22
ESM-1b 0.11
ESM-2 0.13
KeAP 0.29
GOProteinGNN $\textbf{0.31}^{*}$
Table 4. Results on remote homology detection task. The evaluation metric is accuracy. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
6. Ablation Tests
We conducted ablation tests on all tasks to comprehensively assess the performance of our proposed algorithm. Consistent trends were observed across tasks; hence, we focus on the results for contact prediction and PPI tasks, which are widely studied benchmarks.
6.1. GKI Ablation
The ablation test for the GKI component is presented in Table 5. The full model results highlight the effectiveness of GOProteinGNN’s model components, where the model demonstrated superior performance across all metrics, underscoring its pivotal role. Ignoring relations through the GKI component implies using GCN (Kipf and Welling, 2017) instead of RGCN (Schlichtkrull et al., 2017) and ignoring edge type during the knowledge graph learning. Without relation learning, we observe a slight decrease in results, emphasizing its importance. Moreover, excluding the GKI component (i.e., without knowledge integration) substantially degrades performance, reinforcing the significance of the GNN mechanism for effective knowledge integration. These experiments highlight the contribution of the complete GKI component to GOProteinGNN’s ability to capture essential biological contexts and interactions, leading to performance improvements.
Method Short Range Contact PPI
P@L P@L/2 P@L/5 F1
GOProteinGNN $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{80.24}^{*}$
-- GKI Ignoring Relations 0.41 0.51 0.64 79.14
-- Without GKI 0.30 0.40 0.52 70.94
Table 5. Ablation test for the GKI component on the contact prediction task, and the PPI task (SHS27K). The metrics are precision scores (P@L, P@L/2, and P@L/5) for short-range contact predictions and the F1 score for the PPI task. The best result is highlighted in bold.
6.2. Number of GKI Components Ablation
We wish to experiment with the number of optimal GKI components. Each GKI component adds an additional knowledge graph injection layer to the model. In the case of GOProteinGNN, the number of GKI components (referred to as $k$) is set to 3. In this ablation test, we reduce the number of GKI components to observe their impact. As shown in Table 6, reducing knowledge injections in GOProteinGNN results in a slight performance decrease. This decrease highlights the significance and meaningfulness of the GKI component.
Method Short Range Contact PPI
P@L P@L/2 P@L/5 F1
GOProteinGNN ($k=3$) $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{80.24}^{*}$
GOProteinGNN ($k=2$) 0.43 0.53 0.64 79.28
Table 6. Ablation test for the number of GKI components ($k$) on the contact prediction task, and the PPI task (SHS27K). The metrics are precision scores (P@L, P@L/2, and P@L/5) for short-range contact predictions and the F1 score for the PPI task. The best result is highlighted in bold.
7. Conclusions
In this paper, we introduced GOProteinGNN, a novel architecture that enhances PLMs by integrating protein knowledge graph information during the creation of amino acid-level representations. Specifically, GOProteinGNN proposed a GKI layer that leverages the [CLS] token representation of the PLM during the encoder stage in a novel manner. Our methodology enables the incorporation of information at both the individual amino acid and protein levels, fostering a comprehensive learning process through graph-based techniques. Furthermore, GOProteinGNN uniquely learns the entire protein knowledge graph during training, allowing it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work. GOProteinGNN showcases its potential across diverse real-world bioinformatics tasks, such as amino acid contact prediction and protein-protein interaction identification. We posit that the observed performance improvements arise from a deeper comprehension of proteins themselves and the associated factual knowledge. This understanding allows generalization for never-before-seen proteins, as demonstrated on downstream tasks. By leveraging both the protein knowledge graph and sequential data modeling, GOProteinGNN effectively propagates information regarding proteins with biological connections. In drug development, it can facilitate virtual screening by prioritizing compounds for binding to specific protein targets, underscoring its potential to provide deeper insights into complex biological phenomena and expediting the drug discovery process.
Appendix A Pre-training Algorithm
@@
Appendix B Pre-training Dataset Description
@@
We employed the ProteinKG25 dataset (Zhang et al., 2022) for pre-training, consistent with all knowledge-enhanced baseline methods, such as OntoProtein (Zhang et al., 2022) and KeAP (Zhou et al., 2023). ProteinKG25 provides a knowledge graph with aligned descriptions and protein sequences, respectively, to GO terms (Ashburner et al., 2000) and protein entities. GO terms can be a molecule, a cellular component, or a biological process. The GO terms and their relations to proteins are described using natural language.
The dataset contains around 5 million triplets, including nearly 600k proteins, 50k attribute terms, and 31 relation terms, with an average of 8.64 relations per protein.
Consistent with KeAP (Zhou et al., 2023), we address the data leakage issue in OntoProtein (Zhang et al., 2022) by removing amino acid sequences present in downstream tasks from the knowledge graph, ensuring they are unseen during pre-training. This inductive setup allows for a more accurate assessment of the model’s generalization to novel proteins. KeAP, after mitigating data leakage, is considered SOTA across all tasks. Given the robust performance demonstrated by KeAP, surpassing methods such as OntoProtein, we focus our comparative analysis on KeAP.
B.0.1. Mitigating Limitations in Knowledge Graphs
Knowledge graphs may contain inconsistencies or incomplete data due to their curation process. To address this, we conducted a thorough analysis to detect inconsistencies within GO relations in our dataset, primarily leveraging lexical approaches and synonym exploitation (Mougin, 2015). Our findings demonstrate that the ProteinKG25 dataset is robust, with inconsistencies accounting for less than 0.02% of the triplets. For a fair comparison, these minor inconsistencies were disregarded. Our primary contribution lies in advancing representation learning through a novel knowledge injection approach, building on prior SOTA methods that have shown promising results with existing knowledge graphs. Future work can explore additional strategies to mitigate these limitations further.
Appendix C Experimental Details
C.1. Task Specific Baselines
Building upon the KeAP (Zhou et al., 2023) baselines, we extend our evaluation to encompass additional SOTA models that are known to perform well on specific tasks within the protein domain. For the contact and homology tasks, we include variants of LSTM (Hochreiter and Schmidhuber, 1997), ResNet (He et al., 2015), and Transformer (Vaswani et al., 2017), as well as ProteinBert (Brandes et al., 2021) for the homology. Following Zhang et al., we examine DPPI (Hashemifar et al., 2018), DNNPPI (Li et al., 2018), PIPR (Chen et al., 2019), and GNN-PPI (Lv et al., 2021) for the protein-protein interactions (PPI) task. Also, we incorporate PIPR (Chen et al., 2019) for affinity estimation and MSA Transformer (Rao et al., 2021) and ProtT5-XL (Elnaggar et al., 2022) for semantic similarity inference.
C.2. Implementation Details
During pre-training, GOProteinGNN is trained for 50k steps using a learning rate of 1e-4, weight decay of 0.02, and a batch size of 512 (proteins) over four GPUs (NVIDIA A40, 48GB Memory each). RGCN layers $L$ are set to 1 and GNN Knowledge Injection (GKI) components (referred to as $k$) are set to 3. Our experiments demonstrated that the single-layer RGCN effectively captures most of the critical information, similar to prior studies. We found that adding additional layers did not yield significant gains and, in some cases, introduced noise into the information. For downstream tasks, we follow the SOTA KeAP (Zhou et al., 2023) to ensure a fair comparison. Specifically, we use hyperparameters from GNN-PPI (Lv et al., 2021) for PPI prediction and PROBE (Unsal et al., 2022) for binding affinity and similarity inference. For TAPE (Rao et al., 2019), we performed a grid search to select the optimal hyperparameters. For full implementation details, refer to the provided code.
C.3. GOProteinGNN Computational Complexity
Efficient handling of computational complexity is crucial for graph-based models such as GOProteinGNN, particularly when applied to large-scale protein knowledge graphs.
In terms of computational complexity, GOProteinGNN, like other GCNs, operates at approximately $O(|V|*|E|)$. Several optimizations, including limiting the number of hops per protein, reduce the complexity to $O(|V|)$, achieving performance akin to KeAP. During application, the complexity of both algorithms is $O(1)$.
In our current setup, the full receptive graph (Wang et al., 2023) for each protein is loaded, enabling effective modeling. For scalability with increasing graph size, studies in graph algorithms offer promising solutions, such as dynamically caching receptive graphs (Wang et al., 2023) for individual nodes. These approaches ensure that our method remains practical for larger datasets.
Appendix D Downstream Tasks Definitions
D.0.1. Contact Prediction